linoleic acid is an essential fatty acid in the human diet, which means that we need to include this acid as part of a healthful diet. linoleic acid is a major constituent of many vegetable oils, including peanut oil, corn oil, and soybean oil. If the mass of a beaker containing 23.0 mL of linoleic acid is 150.07g, and the mass of the empty beaker is 129.32g, what is the density of linoleic acid?


1
Expert's answer
2021-02-01T03:56:48-0500

Density "D= mv"

m=150.07- 129.32=20.75g

V=23cm3

D= 20.75/23

0.9022 g/cm³